# DIP-R1: Deep Inspection and Perception with RL Looking Through and Understanding Complex Scenes

### v1

@@ 5.2.1 Comparisons with Baselines
5.2.2 Analysis
Table 2: Performance comparison with SFT on in-domain CrowdHuman [34] and out-of-domain CityPersons [35], WiderPedestrian [36], and UAVDT [37].
@@ 5.2.2 Analysis
Table 3: The performance comparison (AP50) with SFT at each training step on CrowdHuman.
@@
SFT 16.5 22.6 25.2 30.1 25.9 28.1
DIP-R1 31.3 34.0 34.3 35.4 36.8 37.5
$\Delta$ 14.8 11.4 9.1 5.3 10.9 9.4
Comparison with SFT In this section, we analyze the effectiveness of our framework by comparing it with SFT. First, Table 2 shows SFT performance on in-domain and out-of-domain data. As described in the table, our framework mostly shows superior perception and generalization performance. Notably, SFT methods exhibit particularly poor performance on out-of-domain classes and environments, indicating limited generalization capability beyond the training distribution. We also observe that SFT methods (including other MLLMs), which have high recall values but suffer from significantly low precision, tend to generate a series of meaningless box predictions (false positives); such observations are included in Appendix B. Second, Table 3 presents performance comparison with SFT on CrowdHuman at each training step. For this analysis, we use the same subset of 500 images from CrowdHuman validation set. Our DIP-R1 framework consistently outperforms SFT across all training steps, demonstrating the effectiveness of RL in enhancing visual perception without relying on fully supervised training, using a small amount of data.
Qualitative Analysis We further conduct a qualitative analysis to validate our framework built upon Qwen2.5-VL (7B). Figure 3 shows the visualization results of baseline models and our framework along with ground-truth region boxes. Additionally, Figure 4 illustrates the output of our framework, including the reasoning description, the uncertain regions (green boxes) captured within the observing process, and the residual predictions (orange boxes). As shown in the figure, the reasoning step provides a description of the overall scene context. We also analyze which regions are assigned to the uncertain regions. In the figure, blue boxes represent the uncertain regions, which are included in the observing process, and they tend to capture small or partially visible instances. More qualitative visualization analyses are illustrated in Appendix D
Table 4: Ablation study with Qwen2.5-VL-Instruct (3B) on CrowdHuman.
<think> <look> precision-recall AP50
30.3
34.6
36.6
Ablation Study We also conduct ablation study on CrowdHuman to validate the effectiveness of each component, and the results are shown in Table 4. For this analysis, we also use the same subset of 500 images from CrowdHuman validation set. When there is the reasoning process only, it obtains 30.3 AP50. Then the looking process is additionally adopted together, and it shows another 4.3 gain, achieving 34.6 AP50. Lastly, when a weighted precision-recall is incorporated, it finally obtains 36.6 AP50.
@@
6 Discussion and Limitation
First, as described in Section 4.3, we design a weighted precision-recall reward function, which assigns high rewards when both precision and recall are high. In usual precision-recall curves (which slope downward to the right), there exists an inherent trade-off: high recall is often along with low precision. A simple precision-based reward fails to consider this trade-off. For instance, as shown in Figure 5, it would assign equal rewards for different cases, even though the right prediction is more desirable. It means that there is no any guidance to minimize false negatives and to cover target instances as many as possible. On the other hand, our weighted precision-recall reward modeling distinguishes two cases, encouraging the model to answer in a desirable way covering many target instances. Second, in this work, we explore the effectiveness RL and validity of our designed reward modelings, such as variance-guided look reward, and we focus on evaluating our DIP-R1 framework in the context of fine-grained object detection using extensive in-domain and out-of-domain datasets. However, it can be extended to other tasks, so that we plan to broaden our framework to other multimodal tasks, such as image captioning with multiple region-level groundings, which would further demonstrates its applicability across a wider range of vision-language problems.
7 Conclusion
@@
Appendix A Pseudo-Code for Variance-guided Look Reward
For the details, we provide the pseudo-code for computing variance-guided look reward. Algorithm 1 describes its implementation with required inputs for the better understanding.
Algorithm 1 The Way to Compute Variance-guided Look Reward
2:Initialize set of all boxes $\mathcal{B}\leftarrow\bigcup_{i=1}^{G}\mathcal{B}_{i}$triangleright All predicted boxes from $G$ responses
3:for each box $\mathbf{b}\in\mathcal{B}$ do
4:     $f_{b}\leftarrow 0$
5:     for $i=1,\dots,G$ do
7:              $f_{b}\leftarrow f_{b}+1$triangleright Count the frequency of same box across generations
8:              $continue$
9:         end if
10:     end for
11:     $f_{b}\leftarrow f_{b}/G$
12:     $\text{Var}[\mathbf{k}_{\mathbf{b}}]\leftarrow f_{b}(1-f_{b})$triangleright Compute variance
13:     if $\text{Var}[\mathbf{k}_{\mathbf{b}}]>0$ then
14:         $r_{\mathbf{b}}\leftarrow\alpha\cdot\text{Var}[\mathbf{k}_{\mathbf{b}}]$triangleright Compute box reward based on the variance
15:     else
16:         $r_{\mathbf{b}}\leftarrow-1/\alpha$triangleright Penalty for certain boxes
17:     end if
18:end for
19:$\mathrm{R}_{\text{look}}=[\;]$
21:     Let $\mathcal{B}^{l}_{i}=\{\mathbf{b}_{1},\dots,\mathbf{b}_{N_{L}}\}$
22:     $\mathrm{R}_{i}\leftarrow\frac{1}{N_{L}}\sum_{j=1}^{N_{L}}r_{\mathbf{b}_{j}}$triangleright Compute Look Reward for each response
23:     $\mathrm{R}_{\text{look}}\text{.append}(\mathrm{R}^{i})$
24:end for
25:return $\mathrm{R}_{\text{look}}$
Appendix B Observation of Noisy Box Predictions in SFT
Figure 6: (a) Noisy prediction examples from SFT (Qwen2.5-VL-Instruct 3B) which may lead to high recall but low precision and (b) the corresponding results from our DIP-R1 framework (3B) which performs more accurate prediction.
As mentioned in Section 5.2.2 (Comparison with SFT), we observe that SFT-based methods often achieve high recall but suffer from significantly low precision. Based on the visualization result analysis, we find that SFT models tend to generate a series of meaningless box predictions (i.e., false positives), and it would lead to increase recall and decrease precision. We hypothesize that, this is because SFT models are trained to produce predictions indiscriminately in crowded scenes, aiming to generate as many answers as possible regardless of their correctness. Figure 6 shows several examples. In particular, (a) visualizes the output of the SFT model trained with Qwen2.5-VL-Instruct (3B) [1], which demonstrates this behavior generating many irrelevant predictions regardless of correctness. We speculate that such an overprediction may help minimize the conventional autoregressive loss in crowded scenes. In contrast, as shown in Figure 6(b), our framework generates fewer but more accurate predictions, resulting in more robust performance. Furthermore, Figure 7 shows that DeepSeek-VL2-Small (16B) [23] exhibits a similar overprediction tendency as like SFT.
Figure 7: Noisy prediction examples from DeepSeek-VL2-Small (16B).
Appendix C Instruction Prompts for Baseline Model Evaluation
We provide the details of instruction prompts that we used for evaluating baseline models, Ferret (7B) [40], Qwen2.5-VL-Instruct (3B and 7B) [1], InternVL3 (8B) [70], DeepSeek-VL2-Small (16B) [23], and Groma (7B) [53]. We follow their own default settings as possible as we can. Table 5 shows their instruction prompts. For Qwen series, based on the default instruction prompt “Please output bbox coordinates and names of every {category name} in this image in JSON format.”, we additionally use some phrases, “as many as possible” and “As there might be heavy overlap between $\dots$”. By providing more detailed instruction, these additional phrases improve the performance of baseline Qwen models in crowded scenes.
Table 5: Instruction prompts used for evaluating each baseline models.
Model Instruction Prompt
Ferret [40] Detect all the objects among {person}.
Qwen2.5-VL-Instruct [1] Please output bbox coordinates and names of every person as many as possible in this image in JSON format. As there might be heavy overlap between people, please carefully separate and detect each individual as accurately as possible.
InternVL3 [70] <image> Please detect <ref>person</ref> in the image.
DeepSeek-VL2-Small [23] <image> Locate <|ref|>person<|/ref|> in the given image.
Groma [53] Locate all <p>person</p> in this image.
Appendix D Additional Qualitative Results
Figure 8,  9,  10, and  11 illustrates additional qualitative results of the baseline models and our framework. The images are cropped and enlarged for the better visualization. As described in the figure, the proposed DIP-R1 framework shows robust recognition performance in challenging real-world scenes, recognizing many instances separately and accurately.
Figure 8: The visualization results on CrowdHuman [34].
Figure 9: The visualization results on CityPersons [35].
Figure 10: The visualization results on WiderPedestrian [36].
Figure 11: The visualization results on UAVDT [37].

### v2

@@ V-B Implementation Details
V-C Experimental Result
V-C1 Comparison with Baselines
In this section, we compare the performance of our DIP-R1 framework with various baseline models on four object detection datasets, including in-domain data: CrowdHuman [40] and three out-of-domain data: CityPersons [41], WiderPedestrian [42], and UAVDT [43]. TABLE I shows the performance comparison with the other baseline models: Ferret (7B) [46], Qwen2.5-VL-Instruct (3B, 7B) [38], InternVL3 (8B) [74], DeepSeek-VL2-Small (16B) [25], and Groma (7B) [59]. On CrowdHuman, our framework shows remarkable performance gains over existing baselines, including DeepSeek-VL2-Small (16B) [25] and InternVL3 (8B) [74], which are among the most recently released state-of-the-art MLLMs. Notably, even compared to Groma [59]–a grounding-specialized MLLM which adopts additional region proposal networks, our framework achieves superior performance. On the out-of-domain data, DIP-R1 framework consistently enhances visual perception capabilities across diverse, challenging environments.
V-C2 Qualitative Analysis
We further conduct a qualitative analysis to validate our framework built upon Qwen2.5-VL (7B). Fig. 3 shows the visualization results of baseline models and our framework along with ground-truth region boxes. Additionally, Fig. 4 illustrates the output of our framework, including the reasoning description, the uncertain regions (green boxes) captured within the observing process, and the residual predictions (orange boxes). As shown in the figure, the reasoning step provides a description of the overall scene context. We also analyze which regions are assigned to the uncertain regions. In the figure, blue boxes represent the uncertain regions, which are included in the observing process, and they tend to capture small or partially visible instances. More visualization results are included in Fig. 8.
Figure 5: Qualitative comparative analysis between the SFT baseline and DIP-R1, where both are built upon Qwen2.5-VL-Instruct (3B). As described in (a), SFT usually exhibits noisy prediction results which may lead to high recall but low precision. On the other hand, as shown in (b), DIP-R1 performs more reliable prediction.
TABLE II: Performance comparison with SFT on in-domain (ID) CrowdHuman [40] and out-of-domain (OOD) CityPersons [41], WiderPedestrian [42], and UAVDT [43].
@@ V-C2 Qualitative Analysis
TABLE III: The performance comparison (AP50) with SFT at each training step on CrowdHuman.
@@
SFT 15.9 16.3 21.3 25.5 29.1 33.3
DIP-R1 17.4 25.9 30.8 29.8 33.9 35.5
$\Delta$ 1.5 9.6 9.5 4.3 4.8 2.2
V-C3 Comparison with SFT
In this section, we analyze the effectiveness of our framework by comparing it with SFT. The experiments are conducted with Qwen2.5-VL-Instruct (3B). First, TABLE II shows SFT performance on in-domain and out-of-domain data. As described in the table, our framework mostly shows superior perception and generalization performance. Notably, SFT methods exhibit particularly poor performance on out-of-domain classes and environments, indicating limited generalization capability beyond the training distribution. Second, TABLE III presents performance comparison with SFT on CrowdHuman at each training step. Our DIP-R1 framework consistently outperforms SFT across all training steps, demonstrating the effectiveness of RL in enhancing visual perception without relying on fully supervised training, using a small amount of data. Additionally, Fig. 5 presents a qualitative comparative analysis between the SFT baseline and our framework. As shown in the figure, SFT tends to produce noisy prediction results, which align with its quantitative result of high recall but low precision due to numerous false positives in TABLE II. This sequentially generative behavior may stem from the SFT’s inherent tendency [75, 76] memorizing and imitating frequent patterns in the training data (i.e., crowded consecutive patterns). In contrast, our framework effectively mitigates such unintended behaviors.
V-C4 Ablation Study
We also conduct ablation study on CrowdHuman (CH, in-domain data) and CityPersons (CP, out-of-domain data) to validate the effectiveness of each reward modeling component. TABLE IV shows the experimental result using Qwen2.5-VL-Instruct (3B). When there is the reasoning process only, it obtains 29.5 AP50 on CH and 8.4 AP50 on CP, respectively. Then when the observing process is integrated with the variance-guided look reward, it achieves 33.3 AP50 on CH with 6.3 gain and 11.3 AP50 on CP with 4.8 gain, showing the effectiveness on both in- and out-of-domain data. Lastly, as the weighted precision-recall accuracy reward is incorporated, it obtains 35.5 AP50 on CH and 12.0 AP50 on CP, respectively.
TABLE IV: Ablation study for each reward modeling with Qwen2.5-VL-Instruct (3B) on CrowdHuman (ID) and CityPersons (OOD). The evaluation metric is AP50.
Format Look PR CH CP
<answer> 27.0 8.4
<think>-<answer> 29.5 (+2.5) 9.4 (+1.0)
<think>-<look>-<answer> 33.3 (+6.3) 11.3 (+4.8)
<think>-<look>-<answer> 35.5 (+8.3) 12.0 (+5.5)
@@ V-C4 Ablation Study
VI Discussion
Intuition of Precision-Recall Accuracy Reward As described in Section IV-C, we design a weighted precision-recall reward function, which assigns high rewards when both precision and recall are high. In usual precision-recall curves (which slope downward to the right), there exists an inherent trade-off: high recall is often along with low precision. A simple precision-based reward fails to consider this trade-off. For instance, as shown in Fig. 6, it would assign equal rewards for different cases, even though the right prediction is more desirable. It means that there is no any guidance to minimize false negatives and to cover target instances as many as possible. On the other hand, our weighted precision-recall reward distinguishes two cases, encouraging the model to answer in a desirable way covering many target instances.
Applicability and Purpose of DIP-R1 To verify the general applicability of our framework across diverse vision tasks, we further evaluate DIP-R1 on COCO [77]–a standard benchmark for general object detection–and RefCOCO [78]–a benchmark for visual grounding. The experimental results presented in TABLE V demonstrate the effectiveness of DIP-R1 in extending to other vision-language tasks. Moreover, such analyses confirm that the goal of DIP-R1 is to enhance the fine-grained visual perception capabilities of MLLMs, particularly under visually complex conditions. For instance, our framework can be beneficial for interactive applications such as crowd monitoring and public safety, where it enables identifying individuals performing specific actions or carrying particular objects in crowded environments like airports.
Failure Cases Even though the proposed DIP-R1 obtains noticeable enhancements in visual perception of MLLM under complex scenarios, it still encounters difficulties in recognizing under extremely challenging conditions. Fig. 7 illustrates representative failure cases of our framework. As shown in the top-left and right examples, DIP-R1 struggles to distinguish individual instances when highly repetitive patterns are present. In the bottom-left and right examples, it also fails to accurately perceive small, blurred, or low-light instances that are barely visible.
[…]
Method COCO RefCOCO
DeepSeek-VL2-Tiny (3B) [25] 84.7
Qwen2.5-VL-Instruct (3B) [38] 59.3 85.5
Shikra (7B) [79] 40.3 87.0
DIP-R1 (3B) 66.2 87.5
Figure 7: Failure cases of the proposed framework. The framework still struggles to perceive instances (1) with highly repetitive patterns, (2) which are too small or blurred, and (3) under low-light conditions. The yellow box indicates the perception result, while the blue dotted box is the zoomed-in region of low-light instances.
Limitation In this work, we explore the effectiveness RL and validity of our designed reward modelings, such as variance-guided look reward, and we focus on evaluating our DIP-R1 framework in the context of fine-grained object detection using extensive in-domain and out-of-domain datasets. However, it can be extended to more diverse vision-language tasks, so that we plan to broaden our framework to other tasks, such as image captioning with multiple region-level groundings, which would further demonstrates its applicability across a wider range of vision-language problems.
Figure 8: Additional qualitative visualization results. The results the response results of DIP-R1 including its reasoning descriptions, uncertain regions (green boxes) captured in the observing process, and the prediction results (orange boxes).
VII Conclusion